OpenStudy (anonymous):

I need help with evaluating algibraic experssions(:

OpenStudy (anonymous):

I can help :)

OpenStudy (anonymous):

ok (: thanks the problem is 5/2n+3/3n n=6 (:

OpenStudy (anonymous):

Okay. All you need to do is replace n with 6. So you get: 5/(2*6)+3/(3*6) 5/12+3/18 From here, get a common denominator, which is 36. so (3/3)*(5/12)+(2/2)*(3/18) This gives 15/36+6/36 or 21/36. You can simplify that by dividing top and bottom by 3 giving: 7/12

OpenStudy (anonymous):

Tell me if that doesn't make sense or if you have any other questions :)

OpenStudy (anonymous):

sorry I had to get a pencil for notes(:

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

thank you though. (: it helped can you walk me though one? please

OpenStudy (anonymous):

Of course

OpenStudy (anonymous):

ok n/2 - 1?

OpenStudy (anonymous):

if n=6?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

No problem. So start by substituting 6 in for n. This gives: \[6/2-1\]. 6 divided by 2 is 3. \[3-1\] And 3 minus one is 2. \[2\] So the expression equals two :)

OpenStudy (anonymous):

oh wow thats easy (:

OpenStudy (anonymous):

They all work similarly. But I can do more if you need me to. :)

OpenStudy (anonymous):

ok what about 3 z/5 + 5/z z=10 (:

OpenStudy (anonymous):

Okie dokie. Same as before, start by substituting 10 in for z. This gives: \[(3*10)/5+5/10\] Following the order of operations you do the multiplication first. 3 times 10 is 30. \[30/5+5/10\] From here get a common denominator. In this case it is 10. \[(2/2)*(30/5)+5/10\] \[60/10+5/10\] Then add the numerators giving. 65/10. You can divide by 5 to simplify it a little giving. 13/2. :)

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

You can make it a mixed number or a decimal. As a mixed number it would be 6 1/2. As a decimal that is 6.5.

OpenStudy (anonymous):

I'm not sure which form you would want :P

OpenStudy (anonymous):

haha ok that makes sence I have a big packet with allot of problems and they are all done I just dont have a way to corret them lol.

OpenStudy (anonymous):

does that make sence? lol my teacher isnt in town I have it all done just no way to correct it and I want to pass (:

OpenStudy (anonymous):

Yeah, well give them a try and if you get stuck just post here and I'll try to give you a heading :P

OpenStudy (anonymous):

well I have them done I just need to have them corrected lol. (:

OpenStudy (anonymous):

its a big packet (:

OpenStudy (anonymous):

Julie has 15 new pencils. Eleven of them are priamry colors. what part of the total number of pencils are primary colors.? I got 11/15 (: is that correct

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

ok would you mind going through my packet with me.

OpenStudy (anonymous):

I don't mind at all. But try them first :P

OpenStudy (anonymous):

i have I have them all done I just need them corrected does that make sence?

OpenStudy (anonymous):

Yeah, go for it :P

OpenStudy (anonymous):

ok Ed took a servey of students in Mr. Edwars class. fifteen of a total of twenty one students have pets. what part of the total number of students have pets.? and it is 15/21 (: lol

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok 8/3 as a proper fraction = 2 2/3

OpenStudy (anonymous):

Well, a proper fraction is one such that the numerator is smaller than the denominator. 2 2/3 is a mixed number but I believe that is right.

OpenStudy (anonymous):

ok what about 33 1/2 =67/2 (:

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

3/5 - 7/45 (:

OpenStudy (anonymous):

First, get a common denominator for 45. (9/9)*(3/5)-7/45. 27/45-7/45=20/45. You can then simplify it by dividing it by 5 giving. 4/9

OpenStudy (anonymous):

whoo I got it (:

OpenStudy (anonymous):

19/20-3/4

OpenStudy (anonymous):

Same thing. Find a common denominator of 20. 19/20-(5/5)(3/4)=19/20-15/20=4/20. Dividing by 4 you get 1/5.

OpenStudy (anonymous):

i got it i am just checking my answers in my packet lol (:

OpenStudy (anonymous):

5/8+5/6

OpenStudy (anonymous):

I would use 24 as the common denominator giving 15/24+20/24=35/24.

OpenStudy (anonymous):

ok 34/5x-23/10x = 1 1/2x? (:

OpenStudy (anonymous):

I got 4 1/2x. Getting a common denominator you get 68/10x-23/10x=45/10x=4 1/2x.

OpenStudy (anonymous):

but 3-2 =1?

OpenStudy (anonymous):

is that 3 and 4/5x or 34/5x do you have an x value? Or are you solving for x? Or not?

OpenStudy (anonymous):

it is 3 and 4/5x - 2 and 3/10 x you dont solve for x (:

OpenStudy (anonymous):

Okay, so you have 3 and 8/10x- 2 and 3/10x that gives 1 1/2x. Yes you're right.

OpenStudy (anonymous):

yeah ! (: ok 16 and 5/16- 10 and 7/32 (:

OpenStudy (anonymous):

malevolence19 ?

OpenStudy (anonymous):

Sorry, I was helping someone else. 6 and 3/32 after you get a common denominator.

OpenStudy (anonymous):

its ok what about 5 and 5/8 - 4 and 2/5

OpenStudy (anonymous):

1 and 9/40

OpenStudy (anonymous):

42 and 5/6 + 22 and 7/12 (:

OpenStudy (anonymous):

65 and 5/12

OpenStudy (anonymous):

6 and 3/4x+ 2 and 5/8x (:

OpenStudy (anonymous):

12 and 3/8x

OpenStudy (anonymous):

5 and 7/8 + 6 and 3/5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

12 and 19/40

OpenStudy (anonymous):

4/5-2/3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

15/16+1/2 (:

OpenStudy (anonymous):

1 7/16

OpenStudy (anonymous):

1 and 7/ 16 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

11/12+ 3/4 (: